OpenStudy (anonymous):

find an equation of variation in which y varies inversely as x and the following is true. 1.) y=22, when x=4 2.) y=8, when x=11

OpenStudy (anonymous):

general line equation is y=mx+c plug in given x,y give 2 equations. since y varies inversely as x, the m should be a negative number

OpenStudy (anonymous):

y=88/x y=19/x?